A vector r has a x-component of 2m and a y-component of 3m. Write this vector in polar coordinates.


1
Expert's answer
2021-04-13T06:39:17-0400
"r=\sqrt{2^2+3^2}=3.6\ m\\\theta=\arctan{\frac{3}{2}}=0.9828\ rad\\\vec{r}=(3.6\ m,0.9828\ rad)"